# Sign-Problem-Free Nuclear Quantum Monte Carlo

### v1

@@ Abstract
Quantum Monte Carlo (QMC) methods offer exact solutions for quantum many-body systems but face severe limitations in fermionic systems like atomic nuclei due to the sign problem. While sign-problem-free QMC algorithms exist, they have been confined to simple models with limited predictive power for real nuclei. Here we overcome this barrier by developing a novel lattice nuclear force that is rigorously sign-problem-free for even-even nuclei. This interaction achieves a standard deviation of $\sigma=2.932$ MeV from experimental binding energies for 76 even-even nuclei ($N,Z\leq 28$), matching state-of-the-art phenomenological mean-field models. Key innovations include the first sign-problem-free implementation of spin-orbit coupling for shell evolutions and an efficient QMC-optimized framework for global parameter fitting. Using this approach, we compute binding energies from 4He to 132Sn, symmetric nuclear matter saturation, and reveal novel spin-orbit-driven clustering in light nuclei. This work transforms sign-problem-free QMC into a scalable and predictive nuclear structure tool, establishing a non-perturbative foundation for ab initio calculations extending to heavy nuclei.
@@ Introduction
Notably, for certain symmetry-protected fermionic systems, QMC calculations are immune to the sign problem. In these rare cases, QMC delivers unbiased results with high precision and exceptional scalability. For example, in lattice quantum chromodynamics (QCD), the isospin symmetry between the up and down quarks implies that the fermionic determinant factorizes into a perfect square, guaranteeing positive amplitudes for any gauge-field configuration ProgTheorPhys110-615 . It was further shown that a broader class of interactions without a factorizable determinant can also be sign-problem-free, where positivity is protected by the time-reversal symmetry of the fermion matrix PRB71-155115 ; PRB91-241117 . See Ref. AnnuRevCondMattPhys10-337 and references therein for sign-problem-free QMC applied to phenomenological condensed matter models.
In nuclear physics, however, sign-problem-free QMC algorithms remain underexploited, particularly for ab initio approaches. Such calculations employ complicated high-fidelity nuclear forces derived from fundamental symmetries, unavoidably inducing sign problems PRL111-032501 ; PRL113-192501 ; PRC98-044002 ; Nature630-59 ; PRL132-232502 . Critically, any sign problem, even negligible for light nuclei, grows exponentially with particle number, preventing direct QMC simulations of heavy nuclei. Remedies like perturbation theory PRL128-242501 ; PRR5-L042021 require a nearly sign-problem-free interaction as the non-perturbative zeroth order. Meanwhile, recent ab initio calculations found inconsistencies among few-body, many-body, and nuclear matter predictions PLB736-119 ; PRC104-064312 ; PRC104-054001 ; FewBodySyst64-77 , which stimulates the development of minimal interactions attempting to reconcile these contradictory results PLB797-134863 ; PRL121-072701 ; PRL133-142501 . These minimal interactions typically predict nuclear observables with percent-level accuracy and are potentially solvable using sign-problem-free QMC because of their simplicity. An example is the sign-problem-free Wigner-SU(4) force PLB797-134863 ; PhysRev51-106 ; PRL117-132501 ; PRL127-062501 that successfully reproduces light-nuclei binding energies and charge radii, and describes emergent geometry in 12NatComm14-2777 , clustering in hot nuclear matter PRL125-192502 ; PLB850-138463 , Beryllium isotope spectra PRL134-162503 , and $\alpha$-particle monopole transition PRL132-062501 . These achievements motivate subsequent searches for more accurate sign-problem-free interactions. Here, the core challenge is that time-reversal symmetry enabling sign-problem-free simulations forbids many essential interactions like the tensor force and most spin/isospin-dependent terms PRL95-232502 ; PPNP76-76 ; PRL89-182501 ; PRL98-132501 , severely constraining the available model space.
@@
where the summation spans all lattice sites and the colons mean the normal ordering. Here $\Psi(\bm{n})$ and $\Psi^{\dagger}(\bm{n})$ denote annihilation and creation operators, $M=938.92$ MeV is the nucleon mass, $\nabla^{2}$ is the Laplace operator implemented via fast Fourier transform (FFT), and $C_{2}$, $C_{3}$, $C_{s}$ are coupling constants. We fix the lattice spacing to $a=1.32$ fm throughout and use symbols with overline to denote smeared quantities,
@@
with $s_{\rm NL}$ a real number controlling non-local smearing. Implicit spin/isospin indices in $\Psi(\bm{n})$, $\Psi(\bm{n})^{\dagger}$, and $\sigma_{k}$ are summed. Both local (Eq. (2)) and nonlocal (Eq. (4)) smearing govern the effective interaction range. The relative strength between $s_{\rm L}$ and $s_{\rm NL}$ determines the locality, an essential element for nuclear binding. In particular, the heavy nuclei dissociate into small clusters ($A\leq 4$) at the non-local limit ($s_{\rm L}=0$) and exhibit severe overbinding at the local limit ($s_{\rm NL}=0$PRL117-132501 . The three-body force ($C_{3}$) significantly improves the light nuclei structure PLB797-134863 . The newly introduced spin-orbit term ($C_{s}$) generates requisite shell structure evolution. For $C_{s}=0$ we return to the Wigner-SU(4) interaction thoroughly studied in Ref. PLB797-134863 . We also incorporate a perturbative Coulomb force for protons SM .
@@
where $a_{t}$ is the temporal step and $c(\bm{n})$ is a real auxiliary field. The terms of order $O(C_{s}^{2})$ generated from this transformation are negligible. (Transformation involving 3NF follows SM ). The $c$ fields at each imaginary time step are collectively integrated using the importance sampling Monte Carlo. Without loss of generality, we consider even-even nuclei and take the trial wave function $|\Phi_{T}\rangle$ as a Slater determinant of nucleon wave functions paired by time-reversal symmetry $\mathcal{T}=i\sigma_{y}\mathcal{K}$, where $\mathcal{K}$ is the complex conjugate. These nucleons interact with the $c$-field and propagate independently under Eq. (7). As the single-particle Hamiltonian in Eq. (7) commute with $\mathcal{T}$, the nucleons remain paired during the imaginary-time evolution. The resulting determinant of the fermionic correlation matrix is positively definite, ensuring completely sign-problem-free simulations. (See SM for proof.)
@@ Results
In Fig. 1 we show binding energies for $76$ even-even nuclei with $N,Z\leq 28$. The results show excellent agreement with experimental values, yielding a standard error of $\sigma=2.932$ MeV. For comparison, state-of-the-art density functional theories typically give $\sigma\approx 2$-$3$ MeV in this mass region PRC85-024304 ; mean_field_mass_tables ; PRC104-054312 . Note that our model achieves this level of accuracy with only four parameters trained on six nuclei. This remarkable generalization capability can be attributed to the incorporation of full quantum correlations, with shell effect, pairing, clustering, deformation, collective rotation, and shape fluctuation all treated on an equal footing. Systematic overbinding at large neutron numbers indicates missing components in our nuclear force, such as the pion-exchange potentials dictated by the chiral dynamics RMP81-1773 ; PR503-1 ; frontierinphysics8-98 ; arXiv:2402.14032 .
@@
While existing nuclear QMC approaches using high-fidelity forces remain confined to light/medium-mass systems and require error mitigation, our sign-problem-free method can be directly used to investigate heavy nuclei. In Fig. 2 we plot the QMC results for 100Sn, extrapolated to infinite imaginary time. The statistical errors are sufficiently small to yield total uncertainties below $1$ MeV. We also probe dependencies on temporal step size and box volume, finding only negligible variations. Moreover, The calculated binding energy agrees with experiment within $2$ MeV. These remarkable one-thousandth level computational and theoretical accuracies demonstrate sign-problem-free QMC as a scalable, high-precision approach for heavy nuclei.
@@
4He $-29.0(2)$ $-0.3$ 0.010 $-28.3$
12C $-91.3(1)$ $-13.3$ 0.146 $-92.2$
14C $-104.6(1)$ $-12.7$ 0.121 $-105.3$
16O $-126.9(2)$ $-5.6$ 0.044 $-127.6$
40Ca $-343.0(2)$ $-13.6$ 0.040 $-342.1$
48Ca $-414.5(3)$ $-42.3$ 0.102 $-416.0$
56Ni $-479.3(6)$ $-74.6$ 0.156 $-484.0$
80Zr $-672.1(8)$ $-23.3$ 0.035 $-669.2$
90Zr $-782.1(5)$ $-64.8$ 0.083 $-783.9$
100Sn $-824.7(8)$ $-103.0$ 0.125 $-825.2$
132Sn $-1134.2(27)$ $-110.9$ 0.098 $-1102.8$
Our approach solve the quantum many-body problem non-perturbatively, enabling studies of strong-correlation phenomena like nuclear clustering that emerges from the intricate interplay among shell structure, quantum fluctuations and continuum effects PhysRept432-43 ; RMP90-035004 . Fig. 3 (a) shows the loss function $\chi^{2}$ in Eq. (8) as a function of the spin-orbit coupling constant $C_{s}$, with other parameters fixed to LAT-OPT1 values. A pronounced minimum occurs precisely at the LAT-OPT1 value $C_{s}^{\rm opt}$ (vertical line), confirming that our optimization robustly identifies the physical spin-orbit strength stringently constrained by target nuclei. Fig. 3 (b) displays binding energies for 4He and light $\alpha$-conjugate nuclei (8Be, 12C and 16O) relative to $n$-$\alpha$ breakup thresholds. The 4He binding is independent of $C_{s}$, while 8Be persists unbound against two-$\alpha$ decay. The 3-$\alpha$ (12C) and 4-$\alpha$ (16O) systems are more sensitive to spin-orbit coupling. In the SU(4)-symmetric limit ($C_{s}\to 0)$, 16O is stable, but 12C is unbound against $3\alpha$ dissociation. Conversely, at the other extreme of strong spin-orbit coupling ($C_{s}\gg C_{s}^{\rm opt}$), the 3-$\alpha$ and 4-$\alpha$ binding energies coincide, indicating that 12C becomes deeply bound, while 16O turns unstable against 12C$+\alpha$ breakup. Remarkably, only near the $\chi^{2}$-optimized value $C_{s}^{\rm opt}$ do we reproduce the experimental binding hierarchy. This reveals a direct connection between spin-orbit coupling and effective $\alpha$-cluster interactions, suggesting a unified mechanism governing shell evolution and clustering in light nuclei.
@@
Finally, Fig. 4 shows the predicted equation of state for symmetric nuclear matter. Comparison to the Wigner-SU(4) interaction PLB797-134863 reveals significant improvement. We identify a saturation point at $(\rho_{0},E/A)=(0.175$ fm-3, $-15.4$ MeV)))), consistent with empirical values.
Figure 4: Equation of state for symmetric nuclear matter calculated with LAT-OPT1 (circles) and Wigner-SU(4) PLB797-134863 (diamonds) interactions. Full symbols denote corresponding saturation points. Green rectangle mark the empirical saturation point (0.16±0.01(0.16$\pm$ 0.01( 0.16 ± 0.01 fm-3, $-16.0\pm 1.0$ MeV)))) Bethe1971 .
@@
We have developed a sign-problem-free lattice nuclear force, enabling unbiased QMC simulations across the nuclear chart. This interaction reasonably reproduces experimental binding energies from 4He to 132Sn and nuclear matter saturation with statistical errors of around $1$ MeV at a feasible computational cost. The approach can be directly extended to investigate excited-state spectra PRL106-192501 , density correlations PRL119-222505 , and finite-temperature nuclear properties PRL125-192502 using advanced lattice QMC algorithms. Future refinements will incorporate more nuclei and additional nuclear observables to further constrain the interaction. Moreover, this interaction provides an robust non-perturbative foundation for ab initio calculations with high-fidelity nuclear chiral forces. The missing physics in our model can be incorporated order-by-order perturbatively PRL128-242501 ; Nature630-59 . Finally, this work provides a scalable and systematically improvable computational paradigm that bridges ab initio calculations and phenomenological models.
@@ Acknowledgement
We present the implementation of the three-body force, proof of the positivity, and the details of the Coulomb interaction. Note that these contents can also be found in other literature; we include it for self-sufficiency. We also discuss details of the optimization algorithm, computational complexity, and finite volume effects. We provide results for nucleon-nucleon phase shifts, odd-even staggering in oxygen isotopes, the pure neutron matter equation of state, and charge density distributions. The numerical results and original data are provided as detailed tables and combined figures.
@@
$\langle O\rangle$ $=$ $\lim_{\tau\rightarrow\infty}\frac{\langle\Phi_{T}|e^{-\tau H/2}Oe^{-\tau H/2}|\Phi_{T}\rangle}{\langle\Phi_{T}|e^{-\tau H}|\Phi_{T}\rangle}=\lim_{L_{t}\rightarrow\infty}\frac{\langle\Phi_{T}|M^{L_{t}/2}OM^{L_{t}/2}|\Phi_{T}\rangle}{\langle\Phi_{T}|M^{L_{t}}|\Phi_{T}\rangle}$
@@
Without loss of generality, we assume that $|\Phi_{T}\rangle$ is a Slater determinant composed of single-particle wave functions $\phi_{1},\phi_{2},\ldots,\phi_{A}$. Repeated application of Eq. (16) allows us to rewrite the denominator of Eq. (Proof for the positivity of the amplitude) as
@@
indicating that $v$ and $w$ are orthogonal eigenvectors corresponding to the same real eigenvalue. In this case, the real eigenvalue appears twice in the diagonalized form of $Z$. Consequently, the determinant $\det(Z)$, being the product of all eigenvalues, is always nonnegative.
@@
Gradient descent is an optimization algorithm that leverages gradient information to iteratively update parameters. Given the computational intensity and statistical uncertainties inherent in QMC calculations, complex optimization schemes are infeasible. We adopt a stable and efficient gradient descent method to perform the optimization task, summarized as follows:
@@
    Parameter rescaling: Rescale parameters to ensure comparable variation scales. During early iterations:
    $\theta_{i}\rightarrow\theta_{i}\left|\frac{\partial\chi^{2}}{\partial\theta_{i}}\right|,$
@@
    Gradient computation: At iteration $t$, compute partial derivatives of the loss function. For $C_{2}$, $C_{3}$, and $C_{s}$, directly measure energy contributions from corresponding terms. For $s_{\rm L}$ (nonlinear in binding energy), apply the Feynman-Hellmann theorem for derivative estimation, achieving Monte Carlo uncertainties $<0.1\%$.
@@
Figure 5: optimization of loss function. (Upper) Loss function $\chi^{2}$ versus iteration steps. (Lower) Dimensionless parameters normalized to converged values. Deviations of scaled $s_{\rm L}$ and $C_{2}$ are exaggerated for visibility.
@@
Fig. 6(b) displays statistical uncertainties in binding energies across 16,800 configurations. Uncertainties increase gradually with mass number but remain below 1 MeV for nuclei up to $A\sim 100$. Relative statistical errors (Fig. 6(d)) exhibit rapid convergence in heavy nuclei, stabilizing near one-thousandth for $A\geq 40$.
@@
Figure 9: Binding energies in oxygen isotopes. Binding energy per nucleon $E/A$ versus mass number $A$. Blue circles: even-$A$; red circles: odd-$A$; black crosses: experimental values. Inset: Average phase $\langle e^{i\theta}\rangle$ ($=1$: sign-problem-free).

### v2

@@ Abstract
Quantum Monte Carlo (QMC) methods offer exact solutions for quantum many-body systems but face severe limitations in fermionic systems like atomic nuclei due to the sign problem. While sign-problem-free QMC algorithms exist and provide valuable insights across disciplines, they have been restricted to simple models with limited quantitative predictive power. Here we overcome this barrier by developing a novel lattice nuclear force that is rigorously sign-problem-free for even-even nuclei. This interaction achieves a standard deviation of $\sigma=2.932$ MeV from experimental binding energies for 76 even-even nuclei ($N,Z\leq 28$), matching state-of-the-art phenomenological mean-field models. Key innovations include the first sign-problem-free implementation of spin-orbit coupling for shell evolutions and an efficient QMC-optimized framework for global parameter fitting. Using this approach, we compute binding energies from 4He to 132Sn with unprecedented one-thousandth level numerical precision, reproduce symmetric nuclear matter saturation, and reveal novel spin-orbit-driven clustering in light nuclei. This work transforms sign-problem-free QMC into a scalable and predictive nuclear structure tool, while establishing a high-fidelity, non-perturbative foundation for ab initio calculations of heavy nuclei.
@@ Introduction
Notably, for certain symmetry-protected fermionic systems, QMC calculations are immune to the sign problem. In these rare cases, QMC delivers unbiased results with high numerical precision and exceptional scalability. For example, in lattice quantum chromodynamics (QCD), the isospin symmetry between the up and down quarks implies that the fermionic determinant factorizes into a perfect square, guaranteeing positive amplitudes for any gauge-field configuration ProgTheorPhys110-615 . It was further shown that a broader class of interactions without a factorizable determinant can also be sign-problem-free, where positivity is protected by the time-reversal symmetry of the fermion matrix PRB71-155115 ; PRB91-241117 . See Ref. AnnuRevCondMattPhys10-337 and references therein for sign-problem-free QMC applied to phenomenological condensed matter models.
In nuclear physics, sign-problem-free QMC algorithms are highly desirable but remain underutilized, particularly for nuclear ab initio calculations. Such calculations typically employ complex, high-fidelity nuclear forces derived from fundamental symmetries, which inevitably induce sign problems PRL111-032501 ; PRL113-192501 ; PRC98-044002 ; Nature630-59 ; PRL132-232502 . Crucially, any non-zero sign problem, even negligible in light nuclei, grows exponentially with particle number, preventing direct QMC simulations of heavy nuclei. Remedies like perturbation theory exist PRL128-242501 ; PRR5-L042021 but still require sign-problem-free interactions as the non-perturbative foundation. Consequently, developing sign-problem-free QMC methods would profoundly deepen our first-principles understanding of heavy nuclei. Currently, a prominent sign-problem-free nuclear force is the spin-isospin independent Wigner-SU(4) interaction PLB797-134863 ; PhysRev51-106 ; PRL117-132501 ; PRL127-062501 , which successfully reproduces light-nuclei binding energies and charge radii, describes emergent geometry in 12C NatComm14-2777 , clustering in hot nuclear matter PRL125-192502 ; PLB850-138463 , Beryllium isotope spectra PRL134-162503 , and $\alpha$-particle monopole transitions PRL132-062501 . These significant achievements strongly motivate the search for more accurate sign-problem-free interactions. The core challenge, however, is that the time-reversal principle enabling sign-problem-free simulations inherently forbids essential interactions such as the tensor force and most spin/isospin-dependent terms PRL95-232502 ; PPNP76-76 ; PRL89-182501 ; PRL98-132501 , severely constraining the accessible model space.
@@
where the summation spans all lattice sites and the colons mean the normal ordering. Here $\Psi(\bm{n})$ and $\Psi^{\dagger}(\bm{n})$ denote annihilation and creation operators, $M=938.92$ MeV is the nucleon mass, $\nabla^{2}$ is the Laplace operator implemented via fast Fourier transform (FFT), and $C_{2}$, $C_{3}$, $C_{s}$ are coupling constants. We fix the lattice spacing to $a=1.32$ fm and use natural units $\hbar=c=1$ throughout. We use symbols with overline to denote smeared quantities,
@@
with $s_{\rm NL}$ a real number controlling non-local smearing. Implicit spin/isospin indices in $\Psi(\bm{n})$, $\Psi(\bm{n})^{\dagger}$, and $\sigma_{k}$ are summed. Both local (Eq. (2)) and nonlocal (Eq. (4)) smearing govern the effective interaction range. The relative strength between $s_{\rm L}$ and $s_{\rm NL}$ determines the locality, an essential element for nuclear binding. In particular, the heavy nuclei dissociate into small clusters ($A\leq 4$) at the non-local limit ($s_{\rm L}=0$) and exhibit severe overbinding at the local limit ($s_{\rm NL}=0$) PRL117-132501 . The three-body force ($C_{3}$) significantly improves the light nuclei structure PLB797-134863 . The newly introduced spin-orbit term ($C_{s}$) generates requisite shell structure evolution. For $C_{s}=0$ we return to the Wigner-SU(4) interaction thoroughly studied in Ref. PLB797-134863 . We also incorporate a perturbative Coulomb force for protons SM .
@@
where $a_{t}$ is the temporal step and $c(\bm{n})$ is a real auxiliary field. The terms of order $O(C_{s}^{2})$ generated from this transformation are negligible (Transformation involving 3NF follows SM ). The $c$ fields at each imaginary time step are collectively integrated using the importance sampling Monte Carlo. Without loss of generality, we consider even-even nuclei and take the trial wave function $|\Phi_{T}\rangle$ as a Slater determinant of nucleon wave functions paired by time-reversal symmetry $\mathcal{T}=i\sigma_{y}\mathcal{K}$, where $\mathcal{K}$ is the complex conjugate. These nucleons interact with the $c$-field and propagate independently under Eq. (7). As the single-particle Hamiltonian in Eq. (7) commute with $\mathcal{T}$, the nucleons remain paired during the imaginary-time evolution. The resulting determinant of the fermionic correlation matrix is positively definite, ensuring completely sign-problem-free simulations (See SM for proof). For odd-even and odd-odd nuclei, the nucleons can not be completely paired up and a mild sign problem remains.
@@ Results
In Fig. 1 we show binding energies for $76$ even-even nuclei with $N,Z\leq 28$. The results show excellent agreement with experimental values, yielding a standard error of $\sigma=2.932$ MeV. For comparison, state-of-the-art density functional theories typically give $\sigma\approx 2$-33 MeV in this mass region PRC85-024304 ; mean_field_mass_tables ; PRC104-054312 . Note that our model achieves this level of accuracy with only four parameters trained on six nuclei. This remarkable generalization capability can be attributed to the incorporation of full quantum correlations, with shell effect, pairing, clustering, deformation, collective rotation, and shape fluctuation all treated on an equal footing. Systematic overbinding at large neutron numbers indicates missing components in our nuclear force, such as the pion-exchange potentials dictated by the chiral dynamics RMP81-1773 ; PR503-1 ; frontierinphysics8-98 ; arXiv:2402.14032 .
@@
While existing nuclear QMC approaches using high-fidelity forces remain confined to light/medium-mass systems and require error mitigation, our sign-problem-free method can be directly used to investigate heavy nuclei. In Fig. 2 we plot the QMC results for 100Sn, extrapolated to infinite imaginary time. We find a combined QMC statistical and extrapolation error below $1$ MeV. We also probe dependencies on temporal step size and box volume, finding only negligible variations. Moreover, The extrapolated binding energy reproduces the experiment within $1$ MeV. These unprecedented one-thousandth level numerical precision and remarkable predicative capability demonstrate sign-problem-free QMC as an accurate and scalable approach for heavy nuclei.
@@
4He 28.8-28.8 0.3-0.3 0.010 28.3-28.3
12C 91.3(1)-91.3(1) 13.3-13.3 0.146 92.2-92.2
14C 104.6(1)-104.6(1) 12.7-12.7 0.121 105.3-105.3
16O 126.9(2)-126.9(2) 5.6-5.6 0.044 127.6-127.6
40Ca 343.0(2)-343.0(2) 13.6-13.6 0.040 342.1-342.1
48Ca 414.5(3)-414.5(3) 42.3-42.3 0.102 416.0-416.0
56Ni 479.3(6)-479.3(6) 74.6-74.6 0.156 484.0-484.0
80Zr 672.1(8)-672.1(8) 23.3-23.3 0.035 669.2-669.2
90Zr 782.1(5)-782.1(5) 64.8-64.8 0.083 783.9-783.9
100Sn 824.7(8)-824.7(8) 103.0-103.0 0.125 825.2-825.2
132Sn 1134.2(27)-1134.2(27) 110.9-110.9 0.098 1102.8-1102.8
[…]
@@
Lastly, Fig. 4 shows the predicted equation of state for symmetric nuclear matter. Comparison to the Wigner-SU(4) interaction PLB797-134863 reveals significant improvement. We identify a saturation point at $(\rho_{0},E/A)=(0.175$ fm-3, 15.4-15.4 MeV)), consistent with empirical values.
Figure 4: Equation of state for symmetric nuclear matter calculated with LAT-OPT1 (circles) and Wigner-SU(4) PLB797-134863 (diamonds) interactions. Full symbols denote corresponding saturation points. Green rectangle mark the empirical saturation point $(0.16\pm 0.01$ fm-3, $16.0\pm 1.0$ MeV)) Bethe1971 .
@@
This work presents the first exactly solvable quantum many-body model for atomic nuclei that quantitatively reproduces experimental data across light to heavy nuclei. Exactly solvable quantum many-body problems are rare benchmarks essential for evaluating advanced algorithms, including those based on artificial neural networks RMP94-031003 or quantum computing CPB30-020306 ; EPJConf296-01025 . Here, exact solvability implies achieving the desired precision within polynomial time. For modeling nuclear structure, the proposed sign-problem-free algorithm offers orders of magnitude acceleration compared to calculations using realistic interactions for heavy nuclei. The approach extends directly to investigating excited-state spectra PRL106-192501 , density correlations PRL119-222505 , and finite-temperature nuclear properties PRL125-192502 using advanced lattice QMC algorithms. Furthermore, it establishes a robust non-perturbative foundation for ab initio calculations with complex chiral nuclear forces, enabling perturbative incorporation of missing physics order-by-order PRL128-242501 ; Nature630-59 .
Finally, this work establishes a scalable, systematically improvable computational paradigm that bridges complex ab initio frameworks and simple phenomenological models. Recent ab initio studies have uncovered inconsistencies among few-body systems, many-body nuclei, and nuclear matter predictions PLB736-119 ; PRC104-064312 ; PRC104-054001 ; FewBodySyst64-77 , prompting the development of minimal interactions to resolve these discrepancies PLB797-134863 ; PRL121-072701 ; PRL133-142501 . Our framework represents such a minimal interaction approach, potentially enhancing the explainability of modern ab initio nuclear theory by distilling the essential elements of nuclear binding. This advances the quest for fundamental nuclear interactions and will likely inspire future research across nuclear physics, quantum many-body methods, and related disciplines.
@@ Acknowledgement
Supplemental Materials
Zhong-Wang Niu
Bing-Nan Lu
January 4, 2026
@@
This supplement presents the implementation details of the three-body force, the proof of positivity for our approach, and the treatment of Coulomb interactions. While these elements are documented elsewhere in the literature, they are included here for completeness and self-sufficiency. We further elaborate on: (i) the optimization algorithm employed, (ii) computational complexity analysis, and (iii) finite volume effects. Numerical results and original data are provided in comprehensive tables and composite figures, including nucleon-nucleon phase shifts, odd-even staggering in oxygen isotopes, the pure neutron matter equation of state, and charge density distributions.
@@
$\langle O\rangle$ $=$ $\lim_{\tau\rightarrow\infty}\frac{\langle\Phi_{T}|e^{-\tau H/2}Oe^{-\tau H/2}|\Phi_{T}\rangle}{\langle\Phi_{T}|e^{-\tau H}|\Phi_{T}\rangle}=\lim_{L_{t}\rightarrow\infty}\frac{\langle\Phi_{T}|M^{L_{t}/2}OM^{L_{t}/2}|\Phi_{T}\rangle}{\langle\Phi_{T}|M^{L_{t}}|\Phi_{T}\rangle}$ (14)
@@
Without loss of generality, we assume that $|\Phi_{T}\rangle$ is a Slater determinant composed of single-particle wave functions $\phi_{1},\phi_{2},\ldots,\phi_{A}$. Repeated application of Eq. (16) allows us to rewrite the denominator of Eq. (14) as
@@
indicating that $v$ and $w$ are orthogonal eigenvectors corresponding to the same real eigenvalue. In this case, the real eigenvalue appears twice in the diagonalized form of $Z$. Consequently, the determinant $\det(Z)$, being the product of all eigenvalues, is always nonnegative. We note that for odd-even or odd-odd nuclei the correlation matrix can not be written in the form of Eq. (24) and the resulting determinant could be negative or complex.
Dependence on the lattice spacing
In this work, we adopt a fixed lattice spacing of $a=1.32$ fm, corresponding to a momentum cutoff $\Lambda=\pi/a\approx 471$ MeV. This value is prevalent in recent NLEFT investigations PLB797-134863* ; PRL128-242501* ; Nature630-59* ; PRL132-232502* ; PRL134-162503* . We note that most continuum chiral EFT calculations employ a smooth cutoff $\Lambda$ between $450$ and $550$ MeV NPA747-362* ; PRC68-041001* , yielding similar cutoff dependencies. This cutoff scale also exposes a hidden spin-isospin exchange symmetry originating from QCD PRL127-062501* .
So far, most NLEFT calculations have been performed at a single lattice spacing. Systematic multi-cutoff studies have been limited to nucleon-nucleon scattering EPJA53-83* ; PRC98-044002* ; PRC112-014009* and few-body systems ($A\leq 4$) PLB747-511* ; arXiv:2509.02953* . In this work, we expect that our optimization algorithm also works for other lattice spacings in a reasonable interval and the optimized nuclear force exhibits similar performance, with residual cutoff dependencies that can be estimated within the EFT framework. In typical nuclear EFTs, cutoff-dependencies scale as $O((Q/\Lambda)^{\nu+1})$, where $Q$ is the characteristic momentum and $\nu$ is the order of the EFT interaction RMP92-025004* . Our interaction incorporates an improved leading-order ($O(Q^{0})$) term with partial absorption of next-to-leading-order ($O(Q^{2})$) effects, including finite-range smearing and spin-orbit terms, and next-to-next-to-leading-order ($O(Q^{3})$) three-body forces. We thus make a rough estimate $\nu\approx 2$. In nuclei, $Q\sim\sqrt{ME}\approx 100$ MeV, with $M$ the nucleon mass and $E$ the typical single-nucleon separation energy. For sufficiently large $\Lambda$ (e.g., $\Lambda\geq 300$ MeV, corresponding to $a\leq 2$ fm), cutoff dependencies are suppressed to the order of $O((100/300)^{3})\approx O(0.037)$. These residual lattice effects can be numerically verified and further reduced by including higher-order interactions, which we leave for future work.
Galilean invariance
The interactions introduced in this work explicitly break Galilean invariance, which originates from the essential nonlocality required to reproduce experimental data. In the center-of-mass frame, our local and non-local smearing effectively mimic the $q^{2}$ and $k^{2}$ terms at next-to-leading order in the chiral expansion, while the spin-orbit term reduces to the standard contact coupling $\frac{i}{2}(\bm{q}\times\bm{k})\cdot(\bm{\sigma}_{1}+\bm{\sigma}_{2})$. To our knowledge, no existing nuclear force simultaneously satisfies Galilean invariance, avoids the QMC sign problem, and reproduces experimental binding energies.
@@
Gradient descent is an optimization algorithm that leverages gradient information to iteratively update parameters. The method used in this work is summarized as follows:
@@
    Parameter rescaling: Rescale the parameters to approximately normalize the partial derivatives. During early iterations:
    $\theta_{i}\rightarrow\theta_{i}\left|\frac{\partial\chi^{2}}{\partial\theta_{i}}\right|_{t=0},$
@@
    Gradient computation: For each iteration, the partial derivatives of the loss function $\partial\chi^{2}/\partial\theta_{i}$ must be computed. Fig. 5 compares two methods for computing these partial derivatives.
    One straightforward approach is to directly calculate the energies using two sets of parameters differing by small variations $\theta_{i}$ and $\theta_{i}+\delta\theta_{i}$, and then approximate the derivative via finite difference. However, these two calculations are independent and their combined statistical errors are typically similar in magnitude to the difference being measured. Consequently, the resulting derivative estimate often suffers from large statistical uncertainty.
    Alternatively, we can apply the Feynman-Hellmann theorem (FHT). This allows us to directly measure the derivatives as observables within a single simulation run. Crucially, this approach significantly reduces the statistical errors. The key reason for this improvement is that the FHT method inherently captures the correlation between the energies at the slightly different parameter values within the same ensemble. Furthermore, this method enables the simultaneous measurement of derivatives with respect to any number of parameters.
    This precise and efficient estimation of the derivatives forms the essential foundation of our optimization algorithm. The algorithm’s capability to handle derivatives effectively makes it suitable for application to a wider range of target nuclei and larger sets of adjustable parameters.
    Figure 5: Comparison of methods for computing derivatives. (Upper panel) Ground state energies of 16O calculated with $s_{\rm L}$= 0.0808 (blue circles) and 0.0818 (red diamonds), respectively. Other parameters take the LAT-OPT1 values. Errorbars denote independent statistical errors in each calculation. (Lower panel) Blue circles represent the difference between the two groups of points in the upper panel, with combined errors estimated with the error propagation formula $e=\sqrt{e_{1}^{2}+e_{2}^{2}}$. Red squares denote the results from the Feynman-Hellmann theorem, with the statistical errors smaller than the symbols.
@@
    Figure 6: optimization of loss function. (Upper) Loss function $\chi^{2}$ versus iteration steps. (Lower) Dimensionless parameters normalized to converged values. Deviations of scaled $s_{\rm L}$ and $C_{2}$ are exaggerated for visibility. We divide each parameter by its value at the last iteration in order to present the convergence of all parameters within the same figure in an illustrative manner.
@@
Fig. 7(b) displays statistical uncertainties in binding energies across 16,800 configurations. Uncertainties increase gradually with mass number but remain below 1 MeV for nuclei up to $A\sim 100$. Relative QMC statistical errors (Fig. 7(d)) exhibit rapid convergence in heavy nuclei, stabilizing near one-thousandth for $A\geq 40$.
@@
It is interesting to discuss the possible application of our method to even heavier nuclei, such as 208Pb or actinides. Based on the approximately linear scaling observed in Fig. 7, we estimate that computations for nuclei with $A\approx 200$ will require computational time roughly twice that for 100Sn. This remains computationally feasible with current supercomputing resources.
However, heavier nuclei exhibit larger radii, necessitating a larger simulation box size to mitigate finite-volume effects. The required box size should scale proportionally to the nuclear radius $R$, which itself scales as $A^{1/3}$. Given that our algorithm scales as $O(L^{3})$ with box size, this volume increase introduces an additional factor of $A$ into the overall scaling. Consequently, we estimate the scaling for precision calculations of heavy nuclei to be $O(A^{2})$. This implies that simulating 208Pb would require approximately four times the computational resources as 100Sn.
For such simulations, addressing challenges like efficient memory management and re-optimization of the effective interaction will be essential prerequisites, along with GPU acceleration to facilitate the calculations.
@@
Figure 10: Binding energies in oxygen isotopes. Binding energy per nucleon $E/A$ versus mass number AA. Blue circles: even-AA; red circles: odd-AA; black crosses: experimental values. Inset: Average phase $\langle e^{i\theta}\rangle$ (=1=1: sign-problem-free).